# MoTE: Mixture of Ternary Experts for Memory-efficient Large Multimodal Models

### v1

@@ 4.3 Compatibility with post-training quantization
We apply GPTQ FAHA (22) and AWQ LTT+ (24) at various bit-widths and report the best results given the same expert memory footprint. We use 512 samples with the length of 2048 tokens from Stage III’s data as the calibration set. For MoE-LLaVA, all full-precision experts are quantized, resulting in expert memory footprints of 2.2GB and 4.5GB under INT4 quantization for the 1.5B and 3B models, respectively. To ensure a fair comparison, we quantize the shared expert of MoTE to INT8 using RTN DLBZ (22). Additionally, we extend the comparison to scenarios with lower memory constraints. For expert memory footprints of 1.6GB and 3.4GB in the 1.5B and 3B models, MoE-LLaVA’s experts are quantized to 3-bit integers using GPTQ, while the shared experts of MoTE are quantized to INT4.
@@ 4.4 Scaling with more data
To examine whether our method is friendly for scaling with data, we train a 1.5B MoTE model with more data during ternary up-cycling. We adopt the same data recipe for Stage I and Stage II as shown in Section 4.1. Then we use a full set of MammoTH-VL GZB+ (24) for Stage III, which contains 10 million samples, each associated with a single image. Every dense layer is replaced with an MoTE layer with one full-precision shared expert and four routed ternary experts. The training steps is set as 40k. The other hyper-parameters are consistent with the setup presented in Section 4.1.
Table 4 summarizes the zero-shot accuracy of MoTE and the baselines across various multimodal reasoning and general VQA tasks. For the baselines, we use their reported scores when available; otherwise, we evaluate the open-sourced models using the same prompts as ours to ensure a fair comparison. As shown in Table 4, although MoTE-1.5B is only trained with 21.6B tokens, our model achieves an improvement of 2.0% average accuracy compared to Qwen2-VL-2B WBT+ (24). Furthermore, MoTE outperforms the larger dense model with fewer FLOPs. Specifically, MoTE outperforms MiniCPM-V-2.0-3B and Phi-3-Vision-4B by a gain of 11.1% and 5.3% accuracy on the testmini set of MathVista.
@@ 4.5 Ablation studies
Precision of Routed Expert MMMU (val) MMB (en test) AI2D (test) ChartQA (test) Seed2+ (test) MMS (test) Avg.$\uparrow$
@@
Precision of Shared Expert Initialize from FFN MMMU (val) MMB (en test) AI2D (test) ChartQA (test) Seed2+ (test) MMS (test) Avg.$\uparrow$
@@
Ternary Training Full-Precision Training MMMU (val) MMB (en test) AI2D (test) ChartQA (test) Seed2+ (test) MMS (test) Avg.$\uparrow$
@@
We investigate the impact of expert precision on the performance of MoTE. Specifically, we compare ternary (i.e., 1.58-bit) up-cycling to 1-bit up-cycling with BWN RORF (16) as the weight quantizers. Both models are up-cycled from Qwen2.5-1.5B with SigLIP-L as the vision encoder to ensure a fair comparison. As shown in Table 5, using binary experts results in performance degradation across most tasks. Similar findings have been reported in the quantization-aware training of BERT models BZH+ (21), where transitioning from ternary to binary weights leads to a substantially more complex and irregular loss landscape, making optimization notably more difficult. Above all, ternary up-cycling is a memory-effective and high-performance solution for MoE models.
@@ 5 Analysis
We visualize the routing distribution of all tokens in MoTE-1.5B on the en-test split of the MMBench dataset. As shown in Figure 2(a), expert utilization across all tokens is well-balanced. To further investigate modality-specific behavior, we present the routing distributions for text and image tokens separately in Figures 2(b) and 2(c), respectively. Notably, text and image tokens exhibit distinct routing patterns. For example, expert #1 is frequently activated for image tokens in the first layer and the final five layers. Additional visualizations across various tasks are provided in Appendix D.1. We observe that routing distributions remain largely consistent across different tasks, suggesting that the experts in MoTE specialize based on modality rather than task-specific features. Moreover, we include per-expert routing distributions by modality in Appendix D.2. Interestingly, some experts exhibit clear modality preferences despite the absence of explicit modality conditioning during training. To better understand expert specialization, we further apply PCA Pea (01) to extract the top-10 routing pathways for text and image tokens. More visualizations are included in AppendixD.3. These findings enhance our understanding of MoTE’s behavior and workflow from a token-level perspective.
@@ Appendix B Hyper-parameters
We utilize torch.compile to compile the PyTorch code in the quantization into optimized kernels, which significantly speed up the training of MoTE. As for the training of 1.5B model’s up-cycling in Stage III, MoTE costs 43.3 hours on 16 NVIDIA A100 cards (40GB), while MoE-LLaVA uses 41.8 hours. Above all, MoTE has similar training time compared to full-precision up-cycling MoE-LLaVA.

### v2

@@ 4.3 Compatibility with post-training quantization
We apply GPTQ Frantar et al. (2022) and AWQ Lin et al. (2024b) at various bit-widths and report the best results given the same expert memory footprint. We use 512 samples with the length of 2048 tokens from Stage III’s data as the calibration set. For MoE-LLaVA, all full-precision experts are quantized, resulting in expert memory footprints of 2.2GB and 4.5GB under INT4 quantization for the 1.5B and 3B models, respectively. To ensure a fair comparison, we quantize the shared expert of MoTE to INT8 using RTN Dettmers et al. (2022). Additionally, we extend the comparison to scenarios with lower memory constraints. For expert memory footprints of 1.6GB and 3.4GB in the 1.5B and 3B models, MoE-LLaVA’s experts are quantized to 3-bit integers using GPTQ, while the shared experts of MoTE are quantized to INT4.
@@ 4.4 Scaling with more data
To examine whether our method is friendly for scaling with data, we train a 1.5B MoTE model with more data during ternary up-cycling. We adopt the same data recipe for Stage I and Stage II as shown in Section 4.1. Then we use a full set of MammoTH-VL Guo et al. (2024) for Stage III, which contains 10 million samples, each associated with a single image. Every dense layer is replaced with an MoTE layer with one full-precision shared expert and four routed ternary experts. The training steps is set as 40k. The other hyper-parameters are consistent with the setup presented in Section 4.1.
Table 4 summarizes the zero-shot accuracy of MoTE and the baselines across various multimodal reasoning and general VQA tasks. For the baselines, we use their reported scores when available; otherwise, we evaluate the open-sourced models using the same prompts as ours to ensure a fair comparison. As shown in Table 4, although MoTE-1.5B is only trained with 21.6B tokens, our model achieves an improvement of 2.0% average accuracy compared to Qwen2-VL-2B Wang et al. (2024c). Furthermore, MoTE outperforms the larger dense model with fewer FLOPs. Specifically, MoTE outperforms MiniCPM-V-2.0-3B and Phi-3-Vision-4B by a gain of 11.1% and 5.3% accuracy on the testmini set of MathVista.
@@ 4.5 Ablation studies
[…]
@@
[…]
@@
[…]
@@
We investigate the impact of expert precision on the performance of MoTE. Specifically, we compare ternary (i.e., 1.58-bit) up-cycling to 1-bit up-cycling with BWN Rastegari et al. (2016) as the weight quantizers. Both models are up-cycled from Qwen2.5-1.5B with SigLIP-L as the vision encoder to ensure a fair comparison. As shown in Table 5, using binary experts results in performance degradation across most tasks. Similar findings have been reported in the quantization-aware training of BERT models Bai et al. (2021), where transitioning from ternary to binary weights leads to a substantially more complex and irregular loss landscape, making optimization notably more difficult. Above all, ternary up-cycling is a memory-effective and high-performance solution for MoE models.
@@ 5 Analysis
We visualize the routing distribution of all tokens in MoTE-1.5B on the en-test split of the MMBench dataset. As shown in Figure 2(a), expert utilization across all tokens is well-balanced. To further investigate modality-specific behavior, we present the routing distributions for text and image tokens separately in Figures 2(b) and 2(c), respectively. Notably, text and image tokens exhibit distinct routing patterns. For example, expert #1 is frequently activated for image tokens in the first layer and the final five layers. Additional visualizations across various tasks are provided in Appendix D.1. We observe that routing distributions remain largely consistent across different tasks, suggesting that the experts in MoTE specialize based on modality rather than task-specific features. Moreover, we include per-expert routing distributions by modality in Appendix D.2. Interestingly, some experts exhibit clear modality preferences despite the absence of explicit modality conditioning during training. To better understand expert specialization, we further apply PCA Pearson (1901) to extract the top-10 routing pathways for text and image tokens. More visualizations are included in AppendixD.3. These findings enhance our understanding of MoTE’s behavior and workflow from a token-level perspective.
@@ Appendix B Hyper-parameters
We utilize torch.compile to compile the PyTorch code in the quantization into optimized kernels, which significantly speed up the training of MoTE. Above all, MoTE has similar training time compared to full-precision up-cycling MoE-LLaVA.